b What is the pröbablly at u tu 16 Light bulbs are sold in packs of 10. A quality inspector selects two bulbs at random without replacement. I both bulbs are defective the pack is rejected. If neither are defective the pack is accepted. If one of th defective the inspector selects two more from the bulbs remaining in the pack and rejects the pack if one or both are defective. What are the chances that a pack containing 4 defective bulbs will in fact be accepted? e bulbs is
Answer #1

total bulbs =10

defective bulbs =4

non defective bulbs=6

binomial distribution

P(accepting)=P( both inspected bulbs are non defective)+P(one bulb defective and next two inspected are non defective )

=6/10*5/9 + 4/10*6/9*5/8 = 0.5

so, 50% chance of accepting
Answer #2

P(accepted at first trial)=6/10 x 5/9 = 1/3
Answer #3

P(accepted at second trial)=6/10 x 4/9 x 5/8 x 4/7 x 2 = 4/21


P(accepted) = 1/3 + 4/21 = 11/21